I need four consecutive even integers whose sum is 340. thanks
Answer by venugopalramana(3286)

I need four consecutive even integers whose sum is 340. thanks
ANY INTEGER IS N..EVEN INTEGER IS 2N.....NEXT 3 ARE..2 MORE THAN THE PREVIOUS ONES... 2N+2,2N+4,2N+6
THEIR SUM IS 2N+2N+2+2N+4+2N+6=8N+12=340
8N=340-12=328
N=41
HENCE THE EVEN INTEGERS ARE 2*41=82,84,86,88